OpenStudy (anonymous):

State whether the following sequences are arithmetic; and if so, state the common difference. 48, 43, 38, 33, … 5 -5 not arithmetic 13

OpenStudy (anonymous):

D = t2 - t1

OpenStudy (anonymous):

43 -48 = -5

OpenStudy (anonymous):

d=a2-a1 so 43-48 or 38-43 = -5

OpenStudy (anonymous):

thanks people